OpenStudy (anonymous):

what is the next three terms in the squence 80,20,5,5/4

OpenStudy (mathmate):

80/20=4 20/5=4 5/(5/4) = 4 (5/4) / ? = 4 ....

OpenStudy (amistre64):

yeah, id say keep dividing by 4 :)

OpenStudy (anonymous):

is the next one 1/32?

OpenStudy (mathmate):

hint: \( \frac{5/4}{x} = 4 \) can you find x?